# CLIP-AGIQA: Boosting the Performance of AI-Generated Image Quality Assessment with CLIP

### v2

@@ 4.3 Ablation Studies
The results indicate that removing or changing any single factor leads to a decrease in performance, confirming their contribution to the performance results in Table 3. It is worth noting that CLIP-IQA+ [24] has already validated the importance of learnable context and quality categories, so we only test the impact of regression on CLIP in the quality assessment of generated images. In variant 1, we observed a significant improvement when regression is added. This indicates that the combination of CLIP priors with a simple regression model is already effective.
In variants 2-5, although the impact on the model’s performance is minimal, exploring these variants still provides us with valuable insights to understand and improve CLIP-AGIQA. Variants 2 and 3 are set up similarly to those explored in CoOp [37]. In our investigation of the backbone, we find a similar conclusion: the more advanced the backbone, the better the performance. However, the conclusion from CoOp that having more context tokens leads to better performance is not satisfied when the context length increased from 16 to 32. This can be due to the increased number of parameters making it harder for the model to converge to an appropriate state, warranting further investigation in future work. Additionally, we demonstrate that a “good” initialization does not make much difference, though this is not explicitly included in the table.

### v3

@@ 4.3 Ablation Studies
The results indicate that removing or changing any single factor leads to a decrease in performance, confirming their contribution to the performance results in Table 3. It is worth noting that CLIP-IQA+ [wang2023exploring] has already validated the importance of learnable context and quality categories, so we only test the impact of regression on CLIP in the quality assessment of generated images. In variant 1, we observed a significant improvement when regression is added. This indicates that the combination of CLIP priors with a simple regression model is already effective.
In variants 2-5, although the impact on the model’s performance is minimal, exploring these variants still provides us with valuable insights to understand and improve CLIP-AGIQA. Variants 2 and 3 are set up similarly to those explored in CoOp [zhou2022learning]. In our investigation of the backbone, we find a similar conclusion: the more advanced the backbone, the better the performance. However, the conclusion from CoOp that having more context tokens leads to better performance is not satisfied when the context length increased from 16 to 32. This can be due to the increased number of parameters making it harder for the model to converge to an appropriate state, warranting further investigation in future work. Additionally, we demonstrate that a “good” initialization does not make much difference, though this is not explicitly included in the table.